The sum of the age of a girl and her brother is 26 years.three years age,her age was four times the age of her brother fine their present age..? Plzzzzz help i am having exam
Answer by rothauserc(4718)

let x be the girls age and y be her brothers age, then
x + y = 26
(x-3) = 4(y-3)
solve fist equation for x and substitute in second equation
(26-y-3) = 4y -12
23-y = 4y -12
add y to both sides of =, then add 12 to both sides of =
5y = 35
y = 7 and x = 19
The girl is 19 and her brother is 7